# Deep Koopman Learning of Nonlinear Time-Varying Systems

### v3

@@ Abstract
This paper presents a data-driven approach to approximate the dynamics of a nonlinear time-varying system (NTVS) by a linear time-varying system (LTVS), which is resulted from the Koopman operator and deep neural networks. Analysis of the approximation error between states of the NTVS and the resulting LTVS is presented. Simulations on a representative NTVS show that the proposed method achieves small approximation errors, even when the system changes rapidly. Furthermore, simulations in an example of quadcopters demonstrate the computational efficiency of the proposed approach.
@@ 3.1 Key Idea
The objectives of (10) and (11) are to approximate (2) and (3), respectively. Here, $10<w<$ 1 is a constant that combines the objective of minimizing ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$. In simple terms, ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$ measure the simulation errors in the lifted and original coordinates, respectively.
@@
Applying the existing deep Koopman operator methods developed in [10, 11, 12, 13] to achieve the DKR by solving (9) based on each $\mathcal{B}_{\tau}$ available has two shortcomings. First, computing the pseudo-inverse in (17) and (18) repeatedly while solving (9) becomes computationally expensive as $\tau$ increases. Second, $\theta_{\tau}$ must be initialized for each $\mathcal{B}_{\tau}$, which can be challenging in time-varying systems applications. To overcome these two limitations, one can apply the deep Koopman operator method to approximate the unknown NTVS efficiently, and we propose the following method.
@@ 3.2 Algorithm
where $\chi_{\tau}=\begin{bmatrix}{\mathbf{G}}_{\tau}\\ \mathbf{U}_{\tau}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\beta_{\tau}}$, $\lambda_{\tau}=({\mathbf{I}}_{\beta_{\tau+1}}+\chi_{\tau+1}^{T}(\chi_{\tau}\chi_{\tau}^{T})^{-1}\chi_{\tau+1})^{-1}\in\mathbb{R}^{\beta_{\tau+1}\times\beta_{\tau+1}}$, $\bar{\lambda}_{\tau}=({\mathbf{I}}_{\beta_{\tau+1}}+{\mathbf{G}}_{\tau+1}^{T}({\mathbf{G}}_{\tau}{\mathbf{G}}_{\tau}^{T})^{-1}{\mathbf{G}}_{\tau+1})^{-1}\in\mathbb{R}^{\beta_{\tau+1}\times\beta_{\tau+1}}$ with ${\mathbf{G}}_{\tau+1}\in\mathbb{R}^{r\times\beta_{\tau+1}},\bar{{\mathbf{G}}}_{\tau+1}\in\mathbb{R}^{r\times\beta_{\tau+1}}$ defined in (14).
The proof of Lemma 1 will be given in [15].
@@
To sum up, we have the following algorithm, which is referred to as deep Koopman learning for time-varying systems (DKTV) in the remainder of this manuscript and its pseudocode is provided in [15].
@@
    Repeat steps $2$-$3$ as the unknown NTVS evolves.
@@ 3.3.1 Preliminaries
We then extend the definition of the Koopman operator from [16] to discrete-time NTVS. Let operator $\mathcal{K}^{(k_{0},k)}:\mathcal{F}\rightarrow\mathcal{F}$ act on functions of state space ϕ()italic-$\phi(\cdot)\in\mathcal{F}$ with $\phi(\cdot):\mathcal{M}\times\mathcal{U}\rightarrow\mathbb{C}$ and defined with two parameters $(k_{0},k)$ as
@@
For brevity, we recall the definitions of $\mathcal{F}$, $\mathcal{M}$, and ϕ()italic-$\phi(\cdot)$ from Definition 1. Let $\nu\in\mathbb{R}$ be the positive measurement on $\mathcal{M}$ and assume that $\mathcal{F}=L_{2}(\nu)$. Given a set of linearly independent $\phi_{i}\in\mathcal{F}$ with $i=1,2,\cdots,r$, and define
[…]
@@ 3.3.2 Analysis
Consider a DNN observable function $g(\cdot,\theta_{\tau}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$, let $\bar{\psi}^{h}=[\bar{\psi}_{1}^{h},\bar{\psi}_{2}^{h},\cdots,\bar{\psi}_{n_{h}}^{h}]^{T}\in\mathbb{R}^{n_{h}}$ be its last hidden layer and $\bar{\psi}^{o}=[\bar{\psi}_{1}^{o},\bar{\psi}_{2}^{o},\cdots,\bar{\psi}_{r}^{o}]^{T}\in\mathbb{R}^{r}$ denotes its output layer, where $\bar{\psi}_{i}^{h}:\mathbb{R}\rightarrow\mathbb{R}$ and $\bar{\psi}_{i}^{o}:\mathbb{R}\rightarrow\mathbb{R}$ are generally nonlinear function chosen by user. Note that here $n_{h}$ and $r$ denote the number of nodes of $\bar{\psi}^{h}$ and $\bar{\psi}^{o}$, respectively.
@@
Proof of Lemma 2 is referred to in [15].
@@
Proof of Lemma 3 is given in [15].
@@
Lemma 3 shows the convergence condition of the existing DKO method regarding the DNN observable function $g(\cdot,\theta_{\tau})$ ($n_{h}\rightarrow\infty$). It replaces the convergence condition of the EDMD method [17] ($r\rightarrow\infty$). This conclusion is used in the present work to reduce $\beta_{\tau}\geq r+m$ so that one can track the unknown NTVS based on $\mathcal{B}_{\tau}$ with small batch size.
@@
The proof of Theorem 1 is given in [15].
@@
The proof of Corollary 1 is given in [15].
@@ 4 Numerical Simulations
In this section, we test the proposed algorithm in two examples, one is a simple nonlinear time-varying system (NTVS), and the other one is a quadcopter. In both examples, the system state input data is observed with a fixed time interval of $0.1$s, and $\beta_{\tau}\equiv\beta$, where $\beta\geq r+m$ is an arbitrary positive integer, and $\tau=\lceil\frac{k-\beta_{0}}{\beta}\rceil$, where $k\in\mathbb{K}_{\tau}$.
@@ 4.1 A Simple NTVS
Experiment Setup. The DNN observable function ($g(\cdot,\theta):\mathbb{R}^{2}\rightarrow\mathbb{R}^{6}$) is built with a hidden layer ReLu() that contains $32$ nodes and an output layer Relu() consisting of $6$ nodes. The DNN training process is implemented by choosing optimizer Adam [19] with learning rate = 1e-3 and weight decay rate = 1e-4. Then $\mathcal{B}_{\tau}$ is observed with $\beta=10$ starting from $x_{0}=[1,0]^{T}$.
@@ 4.2 System States Prediction for Single Quadcopter
Experiment Setup. In order to compare the performance of the proposed algorithm with that of the single neural network method, we let $x_{k}\in\mathbb{R}^{12}$ denote the true state of the system, $\bar{x}_{k}\in\mathbb{R}^{12}$ and $\hat{x}_{k}\in\mathbb{R}^{12}$ denote the estimated states of the single DNN and the proposed method, respectively, and $\bar{e}_{k}=\parallel\bar{x}_{k}-x_{k}\parallel$ and $e_{k}=\parallel\hat{x}_{k}-x_{k}\parallel$ denote the estimation errors of the single DNN and DKTV methods, respectively. The data batch $\mathcal{B}_{\tau}$ is collected with $\beta=30$ manually controlling the quadcopter from position $(0,0,0)$ (meter) to $(1,2,3)$ (meter) with the time-varying disturbance force $w_{t}\in\mathbb{R}^{3}$ applied on the quadcopter. $w_{t}$ is generated from the standard normal distribution. For the DKTV method, the DNN observable function $g(\cdot,\theta):\mathbb{R}^{12}\rightarrow\mathbb{R}^{16}$ is built with one hidden layer Gaussian function containing $64$ nodes and one output layer Relu() consisting of 16 nodes. For the single DNN method, DNN $N(\cdot,\nu):\mathbb{R}^{16}\rightarrow\mathbb{R}^{12}$ is built with the same structure as $g(\cdot,\theta)$. Then the DNN training process of both methods is implemented by choosing optimizer Adam [19] with learning rate = 1e-3 and weight decay rate = 1e-4. The optimal vector $\nu\in\mathbb{R}^{p}$ is found by
@@
4.3 Optimal Control based on the Proposed Deep Koopman Learning
In this subsection, we design the optimal controller based on the proposed method for a classic cartpole example from [22], where we make the coefficient of friction of cart on track ($\mu_{c}(t)$) time-varying during the simulation. And we choose the well-studied model predictive control (MPC) method to demonstrate the optimal control application based on the proposed algorithm, of which the optimization problem is formulated as
{mini}
@@ 4.3 Optimal Control based on the Proposed Deep Koopman Learning
$J(u_{i},x_{i})=$ $\sum_{i=k_{\tau}}^{k_{\tau}+l-1}(\bar{g}(x_{i},\theta_{\tau})^{T}\tilde{Q}_{\tau}\bar{g}(x_{i},\theta_{\tau})+u_{i}^{T}Ru_{i})+g(x_{l},\theta_{\tau})^{T}{\tilde{Q}_{l}}g(x_{l},\theta_{\tau}),$
and $l\leq\beta_{\tau}$ is the time horizon; $i$ is the time-index along the defined time horizon; $x_{i}\in\mathbb{R}^{n}$ and $u_{i}\in\mathbb{R}^{m}$ are the system state and control input at time step $i$, respectively; $\bar{g}(x_{i},\theta_{\tau})=g(x_{i},\theta_{\tau})-g(x^{*},\theta_{\tau})$ with $x^{*}$ denoting the goal state; $x_{l}$ is the terminal state; $\mathcal{X}$ denotes the state constraint and $\mathcal{U}$ denotes the control input constraint; $\tilde{Q}_{\tau}=C_{\tau}^{T}QC_{\tau}\in\mathbb{R}^{r\times r}$, $Q\in\mathbb{R}^{n\times n},R\in\mathbb{R}^{m\times m}$ are positive definite matrices.
4.3.1 Time-varying Cartpole Balance
In this example, the physic governing equation of the cartpole is described as
$\ddot{\bar{\theta}}_{t}$ $=\frac{\cos{\theta_{t}}[-F_{t}-ml\dot{\bar{\theta}}_{t}^{2}\sin{\theta_{t}}+\mu_{t}^{c}\text{sgn}{(\dot{x}_{t}})]/(m_{c}+m)}{l[\frac{4}{3}-(m\cos^{2}{\bar{\theta}_{t}})/(m_{c}+m)]}+\frac{g\sin{\theta_{t}}-\mu_{p}\dot{\bar{\theta}}/{ml}}{l[\frac{4}{3}-(m\cos^{2}{\bar{\theta}_{t}})/(m_{c}+m)]},$
For the simulation, we first build the DNN ($g(\cdot,\theta):\mathbb{R}^{4}\rightarrow\mathbb{R}^{6}$) and set $\beta=12$; then the $\mathcal{B}_{\tau}$ is generated by observing the dynamics in (27) with fixed time interval $t_{k_{\tau}+1}-t_{k_{\tau}}\equiv 0.1s$. The approximated dynamics is directly applied to design the MPC controller to keep the cartpole balanced at the up position with $\bar{\theta}=0,\dot{\bar{\theta}}=0$.
Results Analysis: Fig. 7 shows the trajectory of the time-varying cartpole under the MPC control based on the proposed method. From the result, the cartpole is able to keep the desired up position under the DKTV-based MPC control even though $\mu_{k}^{c}$ increases from $0.0005$ to $14.288$.
Figure 7: MPC based on the Approximated Dynamics
Remark 7
In this subsection, we show the MPC design based on approximated linear dynamics from the proposed algorithm. Compared to the existing MPC method for the NTVS proposed in [23, 24, 25], the proposed method requires neither the priori knowledge of the system dynamics nor the distribution information of the system disturbance.
@@ 5 Conclusions
This paper presents a data-driven approach to approximate the dynamics of a nonlinear time-varying system (NTVS) by a linear time-varying system (LTVS), which is resulted from the Koopman operator and deep neural networks. The proposed algorithm is able to approximate well an NTVS by iterative updating its linear approximation as validated by a simple NTVS system in two dimensions. Moreover, controllers developed based on linear approximation perform well in controlling a quadrotor with complex dynamics. Future work includes employing the proposed method to achieve adaptive autonomy in [26, 27] and to serve as robot’s dynamics estimation in human-robot teaming [28, 29].

### v4

@@ Abstract
This paper presents a data-driven approach to approximate the dynamics of a nonlinear time-varying system (NTVS) by a linear time-varying system (LTVS), which results from the Koopman operator and deep neural networks. Analysis of the approximation error between states of the NTVS and the resulting LTVS is presented. Simulations on a representative NTVS show that the proposed method achieves small approximation errors, even when the system changes rapidly. Furthermore, simulations in an example of quadcopters demonstrate the computational efficiency of the proposed approach.
keywords:
Deep neural networks, Koopman operator, nonlinear time-varying systems.
thanks: Email addresses: {hao93, mous}@purdue.edu, {bowen.h, di.wu}@pnnl.gov, wei.pan@manchester.ac.uk.
@@ 3.1 Key Idea
The objectives of (10) and (11) are to approximate (2) and (3), respectively. Here, $0<w<1$ is a constant that combines the objective of minimizing ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$. In simple terms, ${\mathbf{L}}_{1}$ and ${\mathbf{L}}_{2}$ measure the simulation errors in the lifted and original coordinates, respectively.
@@
Applying the existing deep Koopman operator methods developed in [10, 11, 12, 13] to achieve the DKR by solving (9) based on each $\mathcal{B}_{\tau}$ available has two shortcomings. First, computing the pseudo-inverse in (15) and (16) repeatedly while solving (9) becomes computationally expensive as $\tau$ increases. Second, $\theta_{\tau}$ must be initialized for each $\mathcal{B}_{\tau}$, which can be challenging in time-varying systems applications. To overcome these two limitations, such that one can apply the deep Koopman operator method to approximate the unknown NTVS efficiently, we propose the following method.
@@ 3.2 Algorithm
where
$\chi_{\tau}$ $=\begin{bmatrix}{\mathbf{G}}_{\tau}\\ \mathbf{U}_{\tau}\end{bmatrix}\in\mathbb{R}^{(r+m)\times\beta_{\tau}},$
$\lambda_{\tau}$ $=({\mathbf{I}}_{\beta_{\tau+1}}+\chi_{\tau+1}^{T}(\chi_{\tau}\chi_{\tau}^{T})^{-1}\chi_{\tau+1})^{-1}\in\mathbb{R}^{\beta_{\tau+1}\times\beta_{\tau+1}},$
$\bar{\lambda}_{\tau}$ $=({\mathbf{I}}_{\beta_{\tau+1}}+{\mathbf{G}}_{\tau+1}^{T}({\mathbf{G}}_{\tau}{\mathbf{G}}_{\tau}^{T})^{-1}{\mathbf{G}}_{\tau+1})^{-1}\in\mathbb{R}^{\beta_{\tau+1}\times\beta_{\tau+1}}.$
Here, ${\mathbf{G}}_{\tau+1}\in\mathbb{R}^{r\times\beta_{\tau+1}}$ and $\bar{{\mathbf{G}}}_{\tau+1}\in\mathbb{R}^{r\times\beta_{\tau+1}}$ are defined in (14).
The proof of Lemma 1 will be given in Appendix.
@@
To sum up, we have the following algorithm, which is referred to as deep Koopman learning for time-varying systems (DKTV) in the remainder of this manuscript and its pseudocode is provided in Appendix.
@@
    Repeat steps 222-333 as the unknown NTVS evolves.
@@ 3.3.1 Preliminaries
We then extend the definition of the Koopman operator from [15] to discrete-time NTVS. Let operator $\mathcal{K}^{(k_{0},k)}:\mathcal{F}\rightarrow\mathcal{F}$ act on functions of state space $\phi(\cdot)\in\mathcal{F}$ with $\phi(\cdot):\mathcal{M}\times\mathcal{U}\rightarrow\mathbb{C}$ and defined with two parameters $(k_{0},k)$ as
@@
For brevity, we recall the definitions of $\mathcal{F}$, $\mathcal{M}$, and $\phi(\cdot)$ from Definition 1. Let $\nu\in\mathbb{R}$ be the positive measurement on $\mathcal{M}$ and assume that $\mathcal{F}=L_{2}(\nu)$. Given a set of linearly independent $\phi_{i}\in\mathcal{F}$ with $i=1,2,\cdots,r$, and define
$\mathcal{F}_{r}$ $\coloneqq\text{span}\{\phi_{1},\phi_{2},\cdots,\phi_{r}\},$
@@ 3.3.2 Analysis
Consider a DNN observable function $g(\cdot,\theta_{\tau}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{r}$, let
$\bar{\psi}^{h}=[\bar{\psi}_{1}^{h},\bar{\psi}_{2}^{h},\cdots,\bar{\psi}_{n_{h}}^{h}]^{T}\in\mathbb{R}^{n_{h}}$
be its last hidden layer and $\bar{\psi}^{o}=[\bar{\psi}_{1}^{o},\bar{\psi}_{2}^{o},\cdots,\bar{\psi}_{r}^{o}]^{T}\in\mathbb{R}^{r}$ denotes its output layer, where $\bar{\psi}_{i}^{h}:\mathbb{R}\rightarrow\mathbb{R}$ and $\bar{\psi}_{i}^{o}:\mathbb{R}\rightarrow\mathbb{R}$ are generally nonlinear function chosen by user. Note that here $n_{h}$ and $r$ denote the number of nodes of $\bar{\psi}^{h}$ and $\bar{\psi}^{o}$, respectively.
@@
Proof of Lemma 2 is referred to in Appendix.
@@
Proof of Lemma 3 is given in Appendix.
@@
Lemma 3 shows the convergence condition of the existing DKO method regarding the DNN observable function $g(\cdot,\theta_{\tau})$, i.e., $n_{h}\rightarrow\infty$. It replaces the convergence condition of the EDMD method [16], i.e., $r\rightarrow\infty$. This conclusion is used in the present work to reduce $\beta_{\tau}\geq r+m$ so that one can track the unknown NTVS based on $\mathcal{B}_{\tau}$ with small batch size.
@@
The proof of Theorem 1 is given in Appendix.
@@
The proof of Corollary 1 is given in Appendix.
@@ 4 Numerical Simulations
In this section, we first test the proposed algorithm in two examples: one is a simple nonlinear time-varying system (NTVS), and the other is a quadcopter example. In both examples, the system state input data is observed with a fixed time interval of 0.10.10.1s, and $\beta_{\tau}\equiv\beta$, where $\beta\geq r+m$ is an arbitrary positive integer, and $\tau=\lceil\frac{k-\beta_{0}}{\beta}\rceil$, where $k\in\mathbb{K}_{\tau}$. Subsequently, a motivating example is presented to illustrate the effectiveness of the proposed approach in an optimal design setting.
@@ 4.1 A Simple NTVS
Experiment Setup. The DNN observable function ($g(\cdot,\theta):\mathbb{R}^{2}\rightarrow\mathbb{R}^{6}$) is built with a hidden layer ReLu() that contains 323232 nodes and an output layer Relu() consisting of 666 nodes. The DNN training process is implemented by choosing optimizer Adam [18] with learning rate = 1e-3 and weight decay rate = 1e-4. Then $\mathcal{B}_{\tau}$ is observed with $\beta=10$ starting from $x_{0}=[1,0]^{T}$.
@@ 4.2 System States Prediction for Single Quadcopter
Experiment Setup. In order to compare the performance of the proposed algorithm with that of the single neural network method, we let $x_{k}\in\mathbb{R}^{12}$ denote the true state of the system, $\bar{x}_{k}\in\mathbb{R}^{12}$ and $\hat{x}_{k}\in\mathbb{R}^{12}$ denote the estimated states of the single DNN and the proposed method, respectively, and $\bar{e}_{k}=\parallel\bar{x}_{k}-x_{k}\parallel$ and $e_{k}=\parallel\hat{x}_{k}-x_{k}\parallel$ denote the estimation errors of the single DNN and DKTV methods, respectively. The data batch $\mathcal{B}_{\tau}$ is collected with $\beta=30$ manually controlling the quadcopter from position (0,0,0)(0,0,0)( 0 , 0 , 0 ) (meter) to (1,2,3)(1,2,3)( 1 , 2 , 3 ) (meter) with the time-varying disturbance force $w_{t}\in\mathbb{R}^{3}$ applied on the quadcopter. $w_{t}$ is generated from the standard normal distribution. For the DKTV method, the DNN observable function $g(\cdot,\theta):\mathbb{R}^{12}\rightarrow\mathbb{R}^{16}$ is built with one hidden layer Gaussian function containing 646464 nodes and one output layer Relu() consisting of 16 nodes. For the single DNN method, DNN $N(\cdot,\nu):\mathbb{R}^{16}\rightarrow\mathbb{R}^{12}$ is built with the same structure as $g(\cdot,\theta)$. Then the DNN training process of both methods is implemented by choosing optimizer Adam [18] with learning rate = 1e-3 and weight decay rate = 1e-4. The optimal vector $\nu\in\mathbb{R}^{p}$ is found by
@@
4.3 Optimal Control for Time-varying Cartpole
In this subsection, we apply the proposed method to design an optimal controller for the classical cartpole system, as formulated in [21]. The system dynamics are described as follows:
$\ddot{\bar{\theta}}_{t}$ $=\frac{\cos{\theta_{t}}[-F_{t}-ml\dot{\bar{\theta}}_{t}^{2}\sin{\theta_{t}}+\mu_{t}^{c}\text{sgn}{(\dot{x}_{t}})]/(m_{c}+m)}{l[\frac{4}{3}-(m\cos^{2}{\bar{\theta}_{t}})/(m_{c}+m)]}$
$+\frac{g\sin{\theta_{t}}-\mu_{p}\dot{\bar{\theta}}/{ml}}{l[\frac{4}{3}-(m\cos^{2}{\bar{\theta}_{t}})/(m_{c}+m)]},$
@@ 4.3 Optimal Control for Time-varying Cartpole
Results Analysis: Fig. 7 illustrates the trajectory of the time-varying cartpole system regulated by an MPC controller constructed on the learned deep Koopman model, obtained via the proposed approach. The results demonstrate that the controller successfully maintains the cartpole in the upright position, despite significant variation in the friction coefficient $\mu_{k}^{c}$, which increases from 0.00050.00050.0005 to 14.28814.28814.288.
Figure 7: The trajectory generated by the MPC controller using the dynamics learned by DKTV.
@@ 5 Conclusions
This paper presents a data-driven approach to approximate the dynamics of a nonlinear time-varying system (NTVS) by a linear time-varying system (LTVS), which is resulted from the Koopman operator and deep neural networks. The proposed algorithm is able to approximate an NTVS well by iterative updating its linear approximation, as validated by a simple NTVS system in two dimensions. Moreover, controllers developed based on linear approximation perform well in controlling a quadrotor with complex dynamics. Future work includes employing the proposed method to achieve adaptive autonomy in [22] and to serve as robot’s dynamics estimation in human-robot teaming [23, 24].